Find the natural cubic spline interpolant to f(x) = e^(x^3) at the nodes {xi} i=0 = {−2, 0, 2} . Calculate the value of the interpolant at x = 1. What is the error at this point?
Answer #1

Given function is : f(x)=e^{x^3} .

Here, f(-2)=e^{(-2)^3}

i.e., -2)-e

i.e., 20.000335462

and, f(0)=e^{0^3}

i.e., f(0)=e^{0}

i.e., f(0)=1

and, (2)e

i.e., f (2) = е

i.e., f(2) ~ 2980.95798704

DIFFERENCE TABLE :

x f(x) △ f (x) \Delta^2 f(x)
-2 0.000335462
0.999664538
0 1 2978.958322502
2979.95798704
2 2980.95798704

From the Newton's backward formula we have,

a(u + 1 ) * Δ2N4n-2) L * 21 where u = (x-xn)/h.

Here, x = 1, xn = 2, h = 2.

Then, u = (1-2)/2 ,i.e., u = -1/2 ,i.e., u = -0.5

Then we have,

f(1) 2980.95798704(-0.5) 2979.95798704+ 2978,958322502 2!

i.e., 0.25 f (1) 2980.95798704 1489.9789935 * 2978.958322502

i.e., f (1) 1490.978994-372.36979025

i.e., f (1) 1118.60920375

Therefore, the value of the interpolant at x = 1 is 1118.60920375.

Now, putting x = 1 in the function we get,

f(1)=e^{1^3}

i.e., f(1)=e^{1}

i.e., f(1) = e

i.e., f(1) ะ 2.71828 1828

Therefore, the error at this point is = (1118.60920375-2.718281828) = 1115.89092117.